# A Survey of the Self Supervised Learning Mechanisms for Vision Transformers

### v5

@@ Abstract
Vision Transformers (ViTs) have recently demonstrated remarkable performance in computer vision tasks. However, their parameter-intensive nature and reliance on large amounts of data for effective performance have shifted the focus from traditional human-annotated labels to unsupervised learning and pretraining strategies that uncover hidden structures within the data. In response to this challenge, self-supervised learning (SSL) has emerged as a promising paradigm. SSL leverages inherent relationships within the data itself as a form of supervision, eliminating the need for manual labeling and offering a more scalable and resource-efficient alternative for model training. Given these advantages, it is imperative to explore the integration of SSL techniques with ViTs, particularly in scenarios with limited labeled data. Inspired by this evolving trend, this survey aims to systematically review SSL mechanisms tailored for ViTs. We propose a comprehensive taxonomy to classify SSL techniques based on their representations and pre-training tasks. Additionally, we discuss the motivations behind SSL, review prominent pre-training tasks, and highlight advancements and challenges in this field. Furthermore, we conduct a comparative analysis of various SSL methods designed for ViTs, evaluating their strengths, limitations, and applicability to different scenarios.
@@ 1 Introduction
Deep learning-based algorithms have achieved remarkable success across various disciplines, particularly in computer vision (CV) [1, 2] and natural language processing (NLP) [3]. These models often utilize a pretraining task on large datasets to improve performance. This initial step typically serves as an initialization point, after which the model is fine-tuned for specific downstream tasks. Among pre-training strategies for deep learning, self-supervised learning (SSL) has emerged as a prominent approach. SSL approach is driven by two key motivations. Firstly, network trained on extensive datasets learn transferable patterns that reduce overfitting during training. Secondly, parameters derived from large-scale datasets provide effective initialization, allowing faster convergence between different applications [4]. Despite the abundance of unlabeled web data in the era of big data, developing high-quality labeled datasets through human annotation remains expensive. For instance, platforms such as Scale.com [5] charge approximately \$6.4 per image for the labeling task such as image segmentation. Building an image segmentation dataset containing over a million high-quality samples, like JFT-300M, can cost millions of dollars. Usually, this process is time-intensive and inefficient [6]. Moreover, methods of supervised learning can easily learn misleading connections within dataset, which can result in mistakes and thus may be more sensitive to adversarial exploitations. To address these challenges, strategies like active learning [7], semi-supervised learning [8], and SSL [9] have gained traction. Recently, transformers [10] have emerged as an effective neural network (NN) architecture to leverage deeper insights into input data; specifically, ViTs [11] are considered the alternative to CNNs [12, 13, 14, 15, 16, 17] ViTs have been successfully employed in tasks related to computer vision, including object detection [18], recognition [19] and semantic segmentation [20]. In particular, ViTs have demonstrated superior performance in image classification, when trained with large-volume datasets such as JFT-300M [21]. In essence, SSL methods enable Vision Transformers (ViTs) to understand the underlying structures of data without supervision through pretext tasks, the self-designed objectives in a pre-training phase that encourage the model to learn significant patterns within the data itself. The motivation for applying SSL in ViTs is two-fold. Firstly, SSL enables ViTs to utilize large amounts of unlabeled data, fostering robust and generalizable feature representations. This approach is particularly advantageous when labeled data is limited or costly to produce. Secondly, SSL facilitates pre-training ViTs on extensive datasets, which can then be fine-tuned for specific downstream tasks. This process not only improves the performance of the model but also accelerates training convergence [6].
@@
Several interesting surveys related to SSL have been reported. However, these surveys on SSL, serve to particular domains like recommendation systems [22], graphs [23], [24], sequential transfer learning [25], videos [26], algorithms [6], non-sequential tabular data [27] and recent advancements in using deep learning and image-based models for plant disease detection [28]. A recent survey summarize the deep learning models deployed on edge devices, particularly in healthcare and computer vision [29], highlighting the trade-offs between performance and resources usage. Vision Transformers (ViTs) are a prominent area of research in computer vision [30] and have recently demonstrated several breakthroughs. However, the training of ViTs remains a challenge.
@@
Figure 1: Structure of various sections of the survey paper.
@@ 1.1 Survey Structure
Table 1: Table of abbreviations used in this survey, it contains names of models and techniques used for SSL.
@@
SSL Self Supervised Learning
VITs Vision Transformers
@@
CNN Convolutional Neural Networks
NN Neural Networks
GANs Generative Adversarial Networks
VAEs Variational AutoEncoders
BERT Bi-Directional Encoder Representations from Transformers
SimCLR A Simple Framework for Contrastive Learning of Visual Representations
@@
DeiT Data-efficient Image Transformers
@@
CTransPath Contrastive Transformer Pathology
@@
FLSL Feature-Level Self Supervised Learning
MS-COCO Microsoft Common Objects in Context
@@
CIFAR-10 Canadian Institute for Advanced Research
MAE Masked autoencoders
BEIT Bidirectional Encoder representation
IRs Intermediate Representations
@@
HSL Hybrid SSL Framework
@@ 2 Evolution of SSL
In computer vision, SSL methods are generally segmented into contrastive, generative, and predictive approaches. Contrastive methods, such as MoCo [31] and SimCLR [32], aim to learn patterns by contrasting positive and negative samples. Generative methods, such as GANs [33] and VAEs [34], focus on learning representations by generating samples from the learned representation space. Predictive methods, such as BERT [35] and T5 [36], learn representations by predicting some parts of the input data.
One SSL category is instance discrimination, in which the motto of learning is basically focused on learning patterns by differentiating each sample image from others. This method is rather challenging for large volume of datasets. Examples are SimCLR, SwAV, BYOL, MoCO [32, 37, 38, 31].
@@ 2.1 Evolution of ViTs
In the domain of computer vision, ViTs is newly arising development motivated by the success of transformers architecture which is presented in seminal paper ”Attention Is All You Need” [10]. Transformers has revolutionized the learning process by enabling models to learn patterns from substantial volume of unlabeled data. This approach enables global context modeling, distinguishing ViTs from traditional CNNs. Early ViT models demonstrated strong performance in image classification but faced challenges in computational efficiency and data requirements [39].
In computer vision, the traditional approach in image processing tasks is the utilization of convolutional neural networks (CNNs). However, the CNNs have limitations in capturing dependencies across long spatial distances and global context, both of which are crucial for some vision tasks. ViTs address these limitations via self-attention mechanisms to learn and understand the global context and also dependencies over long spatial distances in the input data. ViTs have also shown strong results in image classification, object detection, medical imaging, and image restoration, often surpassing CNNs when trained on large datasets [40].
The initial vision transformer architecture ViT was presented in the paper ”An Image Is Worth 16x16 Words” [11] Since then, various variants of ViTs are proposed, including DeiT [41], Swin Transformer [42], and TNT [43]. These architectures have obtained state-of-the-art performance in different computer vision based tasks, including image classification, object segmentation and object detection.
@@ 2.2 Importance of SSL in ViTs
SSL is especially crucial in the context of ViTs, as it empowers models to learn and understand feature representations from extensive unlabeled data. ViTs rely on extensive datasets to learn meaningful and effective representations while SSL offers a solution by enabling pretraining on large-scale datasets, allowing ViTs to acquire robust and transferable feature representations for downstream tasks [44, 45, 46, 47].
@@ 2.3 Evolution of SSL in Computer Vision
SSL technique is used for representation learning where a model is designed to learn the underlying relationship between its inputs. The framework of Energy-Based Models (EBMs) [48] can effectively describe this objective, where the aim is to assign high energy to inputs that are incompatible and low energy to those that are compatible. The foundation for many self-supervised learning techniques applied to images was laid by Geoffrey Hinton and others in the 1980s through the development of autoencoders and the proposal of greedy layer-wise pretraining [49], where layers of a deep neural network is trained sequentially. However, autoencoders were basically used for learning features and reducing dimensionality. They were impactful at that time as they made it possible to train the first “deep” networks. An analogous approach was Restricted Boltzmann Machines (RBMs), that allows layer-wise training and then combine those layers to build deep belief nets [50]. Although these techniques were eventually replaced by simpler initialization strategies and extended training procedures, they played a vital part in the development of early deep learning architectures.
@@
In 2000s, types of unsupervised learning, including sparse coding and K-means, were employed for self-supervised learning (SSL). These techniques aimed to learn image features without labeled data, laying the groundwork for more complex self-supervised methods. The category of Spatial Context prediction gained attention in SSL for the training of CNNs. In this regard, Pathak et al. (2016) [51] proposed inpainting using context prediction as a pretext task for SSL in images, where the model predicts missing parts of an image. Zhang et al. (2016) [52] explored a straightforward method of the colorization task for self-supervised representation learning. They trained a model to colorize grayscale images, which significantly enhances its capability to understand textures, patterns, and object identities. Noroozi and Favaro (2016) [53] introduced a method where the model learns to solve jigsaw puzzles, using the relative positions of image patches as supervision. Gidaris et al. (2018) [54] introduced RotaNet, which employs a pretext task where the model predicts the rotation angle applied to an image, enhancing feature learning for various orientations. Meanwhile, clustering approaches also progressed in SSL. Caron et al. (2018) [55] proposed DeepCluster [55], a method that iteratively clusters image features and trains a neural network to predict the cluster assignment.
@@
Contrastive learning became a significant approach in self-supervised learning during the era of CNNs [31, 32]. The roots of contrastive learning can be found in early work on metric learning and the development of Siamese networks [56] in the 1990s and early 2000s. Siamese Network architecture, introduced by Bromley et al. in 1993 [56], involved training twin networks to lessen the relationship among embeddings of matching pairs and expanding the gap between embeddings of constrasting pairs. The formalization of contrastive loss by Hadsell, Chopra, and LeCun (2006) [57] was a significant step forward.
@@
In the late 2010s, the concept of employing contrastive learning for self-supervised learning gained attraction. This period saw the development of several influential techniques that leveraged contrastive objectives to learn meaningful patterns from unlabeled data. Dosovitskiy (2016) [58] used instance discrimination as a pretext task. Each image was handled as a different class, and the network was trained to differentiate between differnt variations of the same image. Wu (2018). [59] refined the instance discrimination approach using a memory repository to save negative examples for the contrastive loss. Oord (2018). [60] introduced method known as Contrastive Predictive Coding in 2018, which learns patterns by predicting future observations in latent space using contrastive learning. Chen (2020) [32] introduced a framework known as SimCLR for contrastive learning of visual features that trains by contrasting similar and non-similar pairs of data. This significantly advanced the performance of self-supervised methods in computer vision. Over time, the idea of hybrid learning gained strength and in 2020, Caron (2020) [37] introduced SwAV, which combines contrastive learning with clustering by using swapped assignments between different augmentations of an image.
@@
The idea of self-distillation gives a new dimension to SSL. Grill (2020) [38] boosted the idea of SSL by introducing BYOL, a self-supervised method that avoids the need for negative samples (simplifying contrastive learning) by using two networks to predict the outputs of each other. He et al. [31] introduced MoCo, which extends the idea of SimCLR by using a dynamic dictionary and a momentum encoder for contrastive learning. Facebook AI Research (Caron [61]) presented DINO, which uses self-distillation with no labels, leveraging VITs for SSL.
@@
In 2020, the idea of transformers in vision was introduced. Vision transformers are high-capacity models and need a huge data for model tuning and good generalization. For improving the learning ability of the transformers, Dosovitskiy (2020) [11] adapted SSL techniques like masked image modeling and contrastive learning for ViT. Inspired by BERT (2018) [35], Dosovitskiy et al. [11] masked certain segment tokens and replaced them with trained mask tokens. Later the model was further trained for predicting pixel values directly. However, they found unsupervised pre-training technique notably less useful than pre-training with supervised methods. Before ViTs, the same idea of inpainting was employed in CNNs, where portions of an image were masked out and the model was taught to reconstruct them, this approach to pre-training is termed as MIM (Masked Image Modeling). In ViTs, the MIM was framed as a regression task. This involved initially utilizing an autoencoder, which encode image segments to separate tokens, subsequently, the transformer is pre-trained to estimate the distinct token values for masked tokens. At that time, Bert pre-training of image transformers (BEiT) [62] demonstrated substantial enhanced results in image classification and semantic segmentation compared to earlier supervised and self-supervised methods. However, its training process was intricate as it required a robust autoencoder to transform image patches into discrete tokens.
@@
The SSL techniques based on cross-covariance (correlation) analysis was introduced in 2021. This family originates from the Canonical Correlation Analysis (CCA) framework developed by Hotelling in 1992 [63]. The main idea behind utilizing CCA is to determine the correlation among two variables by examining their cross-covariance matrices. SSL methods based on this idea include VICReg [64], Barlow Twins [65], SWAV [37], and W-MSE [66]. The latest of these methods is VICReg that balances three goals: variance, invariance, and covariance. Controlling the variance along each dimension of the representation helps to prevent collapse, ensuring that the representation maintains diversity. Invariance ensures that two views of the same data are encoded similarly, while covariance promotes various dimensions of the representation to learn distinct features. This balanced approach allows VICReg to effectively leverage cross-covariance matrices for improved self-supervised learning.
@@
Recently, the idea of multi-modality gained attention with the introduction of CLIP by OpenAI in 2021 [67]. CLIP learns image and text embeddings simultaneously by matching corresponding image-caption pairs. Similarly, ALIGN [68] uses self-supervised learning approaches to develop a shared feature space for images and their corresponding captions. The concept of MIM pre-training was streamlined by two concurrent works: MAE by He et al. [69] and SimMIM by Xie et al. [70]. These approaches propose simplified algorithms that immediately rebuild masked parts of image patches instead of using separate image pieces obtained from an encoder as done in BEiT [62]. This idea was inspired by masked language modeling in NLP. These straightforward pre-training strategies performs better as compared to BEiT on classifying images, object detection and semantic segmentation (downstream tasks). MIM has recently dominated this domain by obtaining state-of-the-art performance on ViTs. The essence of this technique is to enhance the network’s capability to capture visual context at patch level using a denoising auto-encoding mechanism. Invariant Joint Embedding Predictive Architecture (IJEPA) [71] has emerged as a promising alternative, addressing some of the limitations inherent in other self-supervised learning methods. Unlike traditional self-supervised approaches that depend heavily on hand-crafted data augmentations, IJEPA focuses on invariant feature learning and joint embeddings. IJEPA relies on predicting the patterns of destination blocks within an image from a single context block, using a strategic masking method to guide the learning process.
@@
RotNet [54] Spatial awareness, Object recognition Complex texture understanding
Jigsaw Puzzles [53] Object parts, Spatial layout Temporal dynamics
Patch Position Prediction [37] Spatial hierarchies Color/texture learning
Colorization [52] Textures, Patterns, Object ID Spatial relationships
Inpainting [51] Texture continuity, Object completion Temporal understanding
Frame Prediction [72] Motion patterns, Temporal consistency Static image features
Temporal Order Verification [73] Temporal understanding Static spatial understanding
SimCLR, MoCo, CPC [32, 31, 60] Feature discrimination Temporal, Cross-modal learning
DeepCluster, SwAV [55, 37] Feature grouping, Discrimination Dynamic scene understanding
BYOL, DINO, SimSiam [38, 61, 74] Consistent feature learning High-frequency detail understanding
Barlow Twins, VICReg [65, 64] Unique feature emphasis Temporal feature learning
Audio-Visual Correspondence [75] Cross-modal learning Isolated visual feature learning
Self-Labelling [76] Simultaneous clustering and representation Limited scalability
Local Aggregation [77] Localized feature learning Global contextual understanding
TiCo [78] Transformation invariance, Covariance contrast Isolated feature learning
Contrastive Multiview Coding [79] Multiview feature learning Single modality data processing
Relative Predictive Coding [80] Relative feature prediction Absolute feature understanding

### v6

@@ Abstract
Advances in deep learning are re-defining how visual data is processed and understand by the machines. Vision Transformers (ViTs) have recently demonstrated prominent performance in computer vision related tasks. However, their performance improves with increasing numbers of labeled data, indicating reliance on labeled data. Humanly annotated data are difficult to acquire and thus shifted the focus from traditional annotations to unsupervised learning strategies that learn structures inside the data. In response to this challenge, self-supervised learning (SSL) has emerged as a promising technique. SSL utilize inherent relationships within the data as a form of supervision. This technique can reduce the dependence on manual annotations and offers a more scalable and resource-effective approach to training models. Taking these strengths into account, it is necessary to assess the combination of SSL methods with ViTs, especially in the cases of limited labeled data. Inspired by this evolving trend, this survey aims to systematically review SSL mechanisms tailored for ViTs. We propose a comprehensive taxonomy to classify SSL techniques based on their representations and pre-training tasks. Furthermore, we highlighted the motivations behind the study of SSL, reviewed prominent pre-training tasks, and highlight advancements and challenges in this field. Furthermore, we conduct a comparative analysis of various SSL methods designed for ViTs, evaluating their strengths, limitations, and applicability to different scenarios.
@@ 1 Introduction
Deep learning-based algorithms have achieved remarkable success across various disciplines, particularly in computer vision (Kalluri et al. 2025, Ren 2024, Zhang et al. 2022, Mendes et al. 2025) and natural language processing (Bosma et al. 2025, Min et al. 2023). These models often utilize a pretraining task on large datasets to improve performance. This initial step typically serves as an initialization point, after which the model is fine-tuned for specific downstream tasks. Among pre-training strategies for deep learning, self-supervised learning (SSL) has emerged as a prominent approach. SSL approach is driven by two key motivations. Firstly, network trained on extensive datasets learn transferable patterns that reduce overfitting during training. Secondly, parameters derived from large-scale datasets provide effective initialization, allowing faster convergence between different applications (Rani et al. ). Despite the abundance of unlabeled web data in the era of big data, developing high-quality labeled datasets through human annotation remains expensive. For instance, platforms such as Scale.com (AI 2024) charge approximately \$6.4 per image for the labeling task such as image segmentation. Building an image segmentation dataset containing over a million high-quality samples, like JFT-300M, can cost millions of dollars. Usually, this process is time-intensive and inefficient (Gui et al. 2024). Moreover, methods of supervised learning can easily learn misleading connections within dataset, which can result in mistakes and thus may be more sensitive to adversarial exploitations. To address these challenges, strategies like active learning (Gal et al. 2017), semi-supervised learning (Kingma et al. 2014), and SSL (Oord et al. 2018) have gained traction. Recently, transformers (Vaswani et al. 2017) have emerged as an effective neural network (NN) architecture to leverage deeper insights into input data; specifically, ViTs (Dosovitskiy et al. 2020) are considered the alternative to CNNs (LeCun et al. 1998, Alipour and Gallegos 2025, Dai et al. 2025, Al-Abri et al. 2025, González-Almagro et al. 2025, Hosny and Mohammed 2025) ViTs have been successfully employed in tasks related to computer vision, including object detection (Li et al. 2022), recognition (Ranzato et al. 2021) and semantic segmentation (Thisanke et al. 2023). In particular, ViTs have demonstrated superior performance in image classification, when trained with large-volume datasets such as JFT-300M (Sun et al. 2017). In essence, SSL methods enable Vision Transformers (ViTs) to understand the underlying structures of data without supervision through pretext tasks, the self-designed objectives in a pre-training phase that encourage the model to learn significant patterns within the data itself. The motivation for applying SSL in ViTs is two-fold. Firstly, SSL enables ViTs to utilize large amounts of unlabeled data, fostering robust and generalizable feature representations. This approach is particularly advantageous when labeled data is limited or costly to produce. Secondly, SSL facilitates pre-training ViTs on extensive datasets, which can then be fine-tuned for specific downstream tasks. This process not only improves the performance of the model but also accelerates training convergence (Gui et al. 2024).
@@
Several interesting surveys related to SSL have been reported. However, these surveys on SSL, serve to particular domains like recommendation systems (Huang et al. 2022), graphs (Yu et al. 2023), (Wu et al. 2021), sequential transfer learning (Mao 2020), videos (Schiappa et al. 2023), algorithms (Gui et al. 2024), non-sequential tabular data (Wang et al. 2025) and recent advancements in using deep learning and image-based models for plant disease detection (Upadhyay et al. 2025). A recent survey summarize the deep learning models deployed on edge devices, particularly in healthcare and computer vision (Xu et al. 2025), highlighting the trade-offs between performance and resources usage. Vision Transformers (ViTs) are a prominent area of research in computer vision (Wang et al. 2025) and have recently demonstrated several breakthroughs. However, the training of ViTs remains a challenge.
@@
Figure 1: Hierarchical overview of the paper’s structure on Self-Supervised Learning (SSL) techniques. The root node of SSL Techniques is divided further into five major sections as shown in the figure, namely: Section 3 Pre-Training Tasks (covers spatial context-based, color and texture based, temporal and sequence based, contrastive and clustering, distillation and momentum-based, redundancy reduction, and cross-modal techniques); Section 4 Regularization Techniques With representative methods being BYOL, MoCo, SimCLR, DINO, Barlow Twins, SwAV, TiCo, CMC, and self-labelling approaches being surveyed). Section 5 explains the evaluation Metrics along with benchmarks and loss functions. Section 6 discusses comparitive analysis of SSL mechanisms. Section 7 discuss about advancements and open challenges while Section 8 discusses future directions.
@@ 1.1 Survey Structure
Table 1: Abbreviations used throughout the survey for models and techniques in self-supervised learning (SSL)
@@
SSL Self-Supervised Learning
ViT Vision Transformer
@@
CNN Convolutional Neural Network
NN Neural Network
GAN Generative Adversarial Network
VAE Variational Autoencoder
BERT Bidirectional Encoder Representations from Transformers
SimCLR Simple Framework for Contrastive Learning of Visual Representations
@@
DeiT Data-efficient Image Transformer
@@
CTransPath Contrastive Transformer for Pathology
@@
FLSL Feature-Level Self-Supervised Learning
COCO Common Objects in Context (Microsoft COCO dataset)
@@
CIFAR-10 Canadian Institute for Advanced Research 10-class dataset
MAE Masked Autoencoder
BEiT Bidirectional Encoder Representation from Images
IR Intermediate Representation
@@
HSL Hybrid Self-Supervised Learning Framework
@@ 2 Evolution of SSL
In computer vision, SSL methods are generally segmented into contrastive, generative, and predictive approaches. Contrastive methods, such as MoCo (He et al. 2020) and SimCLR (Chen et al. 2020), aim to learn patterns by contrasting positive and negative samples. Generative methods, such as GANs (Goodfellow et al. 2020) and VAEs (Kingma et al. 2019), focus on learning representations by generating samples from the learned representation space. Predictive methods, such as BERT (Devlin et al. 2018) and T5 (Raffel et al. 2020), learn representations by predicting some parts of the input data.
One SSL category is instance discrimination, in which the motto of learning is basically focused on learning patterns by differentiating each sample image from others. This method is rather challenging for large volume of datasets. Examples are SimCLR, SwAV, BYOL, MoCO (Chen et al. 2020, Caron et al. 2020, Grill et al. 2020, He et al. 2020).
@@ 2.1 Evolution of ViTs
In the domain of computer vision, ViTs is newly arising development motivated by the success of transformers architecture which is presented in seminal paper ”Attention Is All You Need” (Vaswani et al. 2017). Transformers has revolutionized the learning process by enabling models to learn patterns from substantial volume of unlabeled data. This approach enables global context modeling, distinguishing ViTs from traditional CNNs. Early ViT models demonstrated strong performance in image classification but faced challenges in computational efficiency and data requirements (Elharrouss et al. 2025).
In computer vision, the traditional approach in image processing tasks is the utilization of convolutional neural networks (CNNs). However, the CNNs have limitations in capturing dependencies across long spatial distances and global context, both of which are crucial for some vision tasks. ViTs address these limitations via self-attention mechanisms to learn and understand the global context and also dependencies over long spatial distances in the input data. ViTs have also shown strong results in image classification, object detection, medical imaging, and image restoration, often surpassing CNNs when trained on large datasets (Haruna et al. 2025).
The initial vision transformer architecture ViT was presented in the paper ”An Image Is Worth 16x16 Words” (Dosovitskiy et al. 2020). Since then, various variants of ViTs are proposed, including DeiT (Touvron et al. 2021), Swin Transformer (Liu et al. 2021), and TNT (Han et al. 2021). These architectures have obtained state-of-the-art performance in different computer vision based tasks, including image classification, object segmentation and object detection.
@@ 2.2 Importance of SSL in ViTs
SSL is especially crucial in the context of ViTs, as it empowers models to learn and understand feature representations from extensive unlabeled data. ViTs rely on extensive datasets to learn meaningful and effective representations while SSL offers a solution by enabling pretraining on large-scale datasets, allowing ViTs to acquire robust and transferable feature representations for downstream tasks (Ye et al. 2024, Yun et al. 2022, Wang et al. 2022, Nguyen et al. 2024).
@@ 2.3 Evolution of SSL in Computer Vision
SSL technique is used for representation learning where a model is designed to learn the underlying relationship between its inputs. The framework of Energy-Based Models (EBMs) (LeCun et al. 2006) can effectively describe this objective, where the aim is to assign high energy to inputs that are incompatible and low energy to those that are compatible. The foundation for many self-supervised learning techniques applied to images was laid by Geoffrey Hinton and others in the 1980s through the development of autoencoders and the proposal of greedy layer-wise pretraining (Bengio et al. 2006), where layers of a deep neural network is trained sequentially. However, autoencoders were basically used for learning features and reducing dimensionality. They were impactful at that time as they made it possible to train the first “deep” networks. An analogous approach was Restricted Boltzmann Machines (RBMs), that allows layer-wise training and then combine those layers to build deep belief nets (Salakhutdinov and Hinton 2009). Although these techniques were eventually replaced by simpler initialization strategies and extended training procedures, they played a vital part in the development of early deep learning architectures.
@@
In 2000s, types of unsupervised learning, including sparse coding and K-means, were employed for self-supervised learning (SSL). These techniques aimed to learn image features without labeled data, laying the groundwork for more complex self-supervised methods. The category of Spatial Context prediction gained attention in SSL for the training of CNNs. In this regard, Pathak et al. (2016) (Pathak et al. 2016) proposed inpainting using context prediction as a pretext task for SSL in images, where the model predicts missing parts of an image. Zhang et al. (2016) (Zhang et al. 2016) explored a straightforward method of the colorization task for self-supervised representation learning. They trained a model to colorize grayscale images, which significantly enhances its capability to understand textures, patterns, and object identities. Noroozi and Favaro (2016) (Noroozi and Favaro 2016) introduced a method where the model learns to solve jigsaw puzzles, using the relative positions of image patches as supervision. Gidaris et al. (2018) (Gidaris et al. 2018) introduced RotaNet, which employs a pretext task where the model predicts the rotation angle applied to an image, enhancing feature learning for various orientations. Meanwhile, clustering approaches also progressed in SSL. Caron et al. (2018) (Caron et al. 2018) proposed DeepCluster (Caron et al. 2018), a method that iteratively clusters image features and trains a neural network to predict the cluster assignment.
@@
Contrastive learning became a significant approach in self-supervised learning during the era of CNNs (He et al. 2020, Chen et al. 2020). The roots of contrastive learning can be found in early work on metric learning and the development of Siamese networks (Bromley et al. 1993) in the 1990s and early 2000s. Siamese Network architecture, introduced by Bromley et al. in 1993 (Bromley et al. 1993), involved training twin networks to lessen the relationship among embeddings of matching pairs and expanding the gap between embeddings of constrasting pairs. The formalization of contrastive loss by Hadsell, Chopra, and LeCun (2006) (Hadsell et al. 2006) was a significant step forward.
@@
In the late 2010s, the concept of employing contrastive learning for self-supervised learning gained attraction. This period saw the development of several influential techniques that leveraged contrastive objectives to learn meaningful patterns from unlabeled data. Dosovitskiy (2016) (Alexey et al. 2016) used instance discrimination as a pretext task. Each image was handled as a different class, and the network was trained to differentiate between differnt variations of the same image. Wu (2018). (Wu et al. 2018) refined the instance discrimination approach using a memory repository to save negative examples for the contrastive loss. Oord (2018). (Oord et al. 2018) introduced method known as Contrastive Predictive Coding in 2018, which learns patterns by predicting future observations in latent space using contrastive learning. Chen (2020) (Chen et al. 2020) introduced a framework known as SimCLR for contrastive learning of visual features that trains by contrasting similar and non-similar pairs of data. This significantly advanced the performance of self-supervised methods in computer vision. Over time, the idea of hybrid learning gained strength and in 2020, Caron (2020) (Caron et al. 2020) introduced SwAV, which combines contrastive learning with clustering by using swapped assignments between different augmentations of an image.
@@
The idea of self-distillation gives a new dimension to SSL. Grill (2020) (Grill et al. 2020) boosted the idea of SSL by introducing BYOL, a self-supervised method that avoids the need for negative samples (simplifying contrastive learning) by using two networks to predict the outputs of each other. He et al. (He et al. 2020) introduced MoCo, which extends the idea of SimCLR by using a dynamic dictionary and a momentum encoder for contrastive learning. Facebook AI Research (Caron (Caron et al. 2021)) presented DINO, which uses self-distillation with no labels, leveraging VITs for SSL.
@@
In 2020, the idea of transformers in vision was introduced. Vision transformers are high-capacity models and need a huge data for model tuning and good generalization. For improving the learning ability of the transformers, Dosovitskiy (2020) (Dosovitskiy et al. 2020) adapted SSL techniques like masked image modeling and contrastive learning for ViT. Inspired by BERT (2018) (Devlin et al. 2018), Dosovitskiy et al. (Dosovitskiy et al. 2020) masked certain segment tokens and replaced them with trained mask tokens. Later the model was further trained for predicting pixel values directly. However, they found unsupervised pre-training technique notably less useful than pre-training with supervised methods. Before ViTs, the same idea of inpainting was employed in CNNs, where portions of an image were masked out and the model was taught to reconstruct them, this approach to pre-training is termed as MIM (Masked Image Modeling). In ViTs, the MIM was framed as a regression task. This involved initially utilizing an autoencoder, which encode image segments to separate tokens, subsequently, the transformer is pre-trained to estimate the distinct token values for masked tokens. At that time, Bert pre-training of image transformers (BEiT) (Bao et al. ) demonstrated substantial enhanced results in image classification and semantic segmentation compared to earlier supervised and self-supervised methods. However, its training process was intricate as it required a robust autoencoder to transform image patches into discrete tokens.
@@
The SSL techniques based on cross-covariance (correlation) analysis was introduced in 2021. This family originates from the Canonical Correlation Analysis (CCA) framework developed by Hotelling in 1992 (Thompson 2000). The main idea behind utilizing CCA is to determine the correlation among two variables by examining their cross-covariance matrices. SSL methods based on this idea include VICReg (Bardes et al. 2022), Barlow Twins (Zbontar et al. 2021), SWAV (Caron et al. 2020), and W-MSE (Ermolov et al. 2021). The latest of these methods is VICReg that balances three goals: variance, invariance, and covariance. Controlling the variance along each dimension of the representation helps to prevent collapse, ensuring that the representation maintains diversity. Invariance ensures that two views of the same data are encoded similarly, while covariance promotes various dimensions of the representation to learn distinct features. This balanced approach allows VICReg to effectively leverage cross-covariance matrices for improved self-supervised learning.
@@
Recently, the idea of multi-modality gained attention with the introduction of CLIP by OpenAI in 2021 (Radford et al. 2021). CLIP learns image and text embeddings simultaneously by matching corresponding image-caption pairs. Similarly, ALIGN (Jia et al. 2021) uses self-supervised learning approaches to develop a shared feature space for images and their corresponding captions. The concept of MIM pre-training was streamlined by two concurrent works: MAE by He et al. (He et al. 2022) and SimMIM by Xie et al. (Xie et al. 2022). These approaches propose simplified algorithms that immediately rebuild masked parts of image patches instead of using separate image pieces obtained from an encoder as done in BEiT (Bao et al. ). This idea was inspired by masked language modeling in NLP. These straightforward pre-training strategies performs better as compared to BEiT on classifying images, object detection and semantic segmentation (downstream tasks). MIM has recently dominated this domain by obtaining state-of-the-art performance on ViTs. The essence of this technique is to enhance the network’s capability to capture visual context at patch level using a denoising auto-encoding mechanism. Invariant Joint Embedding Predictive Architecture (IJEPA) (Assran et al. 2023) has emerged as a promising alternative, addressing some of the limitations inherent in other self-supervised learning methods. Unlike traditional self-supervised approaches that depend heavily on hand-crafted data augmentations, IJEPA focuses on invariant feature learning and joint embeddings. IJEPA relies on predicting the patterns of destination blocks within an image from a single context block, using a strategic masking method to guide the learning process.
@@
RotNet (Gidaris et al. 2018) Spatial awareness, Object recognition Complex texture understanding
Jigsaw Puzzles (Noroozi and Favaro 2016) Object parts, Spatial layout Temporal dynamics
Patch Position Prediction (Caron et al. 2020) Spatial hierarchies Color/texture learning
Colorization (Zhang et al. 2016) Textures, Patterns, Object ID Spatial relationships
Inpainting (Pathak et al. 2016) Texture continuity, Object completion Temporal understanding
Frame Prediction (Villegas et al. 2017) Motion patterns, Temporal consistency Static image features
Temporal Order Verification (Misra et al. 2016) Temporal understanding Static spatial understanding
SimCLR, MoCo, CPC (Chen et al. 2020, He et al. 2020, Oord et al. 2018) Feature discrimination Temporal, Cross-modal learning
DeepCluster, SwAV (Caron et al. 2018; 2020) Feature grouping, Discrimination Dynamic scene understanding
BYOL, DINO, SimSiam (Grill et al. 2020, Caron et al. 2021, Chen and He 2021) Consistent feature learning High-frequency detail understanding
Barlow Twins, VICReg (Zbontar et al. 2021, Bardes et al. 2022) Unique feature emphasis Temporal feature learning
Audio-Visual Correspondence (Arandjelović and Zisserman 2017) Cross-modal learning Isolated visual feature learning
Self-Labelling (Asano et al. 2019) Simultaneous clustering and representation Limited scalability
Local Aggregation (Zhuang et al. 2019) Localized feature learning Global contextual understanding
TiCo (Zhu et al. 2022) Transformation invariance, Covariance contrast Isolated feature learning
Contrastive Multiview Coding (Tian et al. 2020) Multiview feature learning Single modality data processing
Relative Predictive Coding (Doersch and Zisserman 2017) Relative feature prediction Absolute feature understanding